# Differentiable Cyclic Causal Discovery Under Unmeasured Confounders

### v1

@@ 5 Discussion
The appendix is structured as follows: Appendix A presents the theoretical foundations of differentiable cyclic causal discovery in the presence of unmeasured confounders, including the proof of Theorem 3.1, and a characterization of the equivalence class of DMGs that maximize the score function. Appendix B provides implementation details of DCCD-CONF and the baselines. Finally, Appendix C provides additional experimental results comparing DCCD-CONF with the baselines.
@@ Appendix A Theory
In this section, we establish the theoretical foundations of differentiable cyclic causal discovery in the presence of unmeasured confounders. We begin by reviewing key definitions and results from prior work that are essential for proving Theorem 3.1, starting with fundamental graph terminology.
@@ A.1 Preliminaries
where $\mathop{\perp}_{p_{\mathcal{G}}}$ denotes conditional independence, and $p_{\mathcal{G}}$ denotes the observational distribution. This is known as the directed global Markov property of $\mathcal{G}$ [49]. However, in general, cyclic graphs do not obey the directed global Markov property as shown by the counterexample below taken from [36, 50].
@@
Forré and Mooij [49] introduced $\sigma$-separation as a generalization of d-separation to extend the directed global Markov property to cyclic graphs. This concept was motivated by applying $d$-separation to the acyclified version of the DMG. Before delving into $\sigma$-separation, we first define the acyclification procedure of a directed mixed graph, following [36].
@@
Note that $\sigma$-separation reduces to $d$-separation for acyclic graphs, that is, when $\mathrm{sc}_{\mathcal{G}}(i)=\{i\}$ for all $i\in\mathcal{V}$. The following result in [49] relates $\sigma$-separation and $d$-separation.
Proposition A.6 ([49]).
@@
Definition A.7 (General directed global Markov property [49]).
@@ A.2 Joint Causal Modelling and Markov properties
In order to incorporate multiple interventional settings into a single causal modeling framework, we follow the joint causal model introduced by [29], where we augment the system with a set of context variables $\bm{C}^{\mathcal{I}}=(\bm{C}_{1},\ldots,\bm{C}_{K})$ each corresponding to a non-empty interventional setting. In this case, $\bm{C}_{k}=\emptyset$ for all $k=1,\dots,K$ corresponds to the observational setting. We construct an augmented graph, denoted by $\mathcal{G}^{\mathcal{I}}$ consisting of both the system variables $\bm{X}$ and the context variables $\bm{C}^{\mathcal{I}}$, such that the $\mathrm{ch}_{\mathcal{G}}(\bm{C}_{k})=I_{k}$, and no context variable has any parent or a spouse. Figure 5 illustrates the augmented graph for the graph from Example A.3 and the intervention sets $\mathcal{I}=\{\emptyset,\{X_{3}\},\{X_{4}\}\}$. The new system containing both the observed variables and the context variables is called the meta system. Finally, given a set of interventional targets $I_{k}\subseteq\mathcal{V}$ and the corresponding context variable $\bm{C}_{k}$, the structural equations governing the observations $\bm{X}$ has the following form:
$\tilde{F}_{i}(\bm{X}_{\mathrm{pa}_{\mathcal{G}}(i)},\bm{C}^{\mathcal{I}}_{\mathrm{pa}_{\mathcal{G}}^{\mathcal{I}}(i)},Z_{i})=\begin{cases}(\bm{C}_{k})_{i},&\text{if }X_{i}\in I_{k},\\ F_{i}(\bm{X}_{\mathrm{pa}_{\mathcal{G}}(i)},Z_{i}),&\text{if }X_{i}\notin I_{k}.\end{cases}$
@@
For an DMG $\mathcal{G}$ and a choice of $\mathbf{F}:\mathbb{R}^{2d}\to\mathbb{R}^{d}$ such that $\mathbf{f}_{x}$ is unique and invertible and $\mathbf{\Sigma}\succ 0$, the structural equations are uniquely solvable with respect to each strongly connected component of $\mathcal{G}$. Morover, the addition of context variables in the augmented graph does not introduce any new cycles. Therefore the meta system forms a simple SCM. Thus, from Theorem A.21 in [36], the distribution $p_{\mathcal{G}^{\mathcal{I}}}$ is unique and it satisfies the general directed global Markov property. ∎
@@
From Proposition A.6, for a DMG $\mathcal{G}$ and a family of interventional targets $\mathcal{I}=\{I_{k}\}_{k=0}^{K}$, any $\sigma$-separation statement in $\mathcal{G}^{\mathcal{I}}$ translates to a $d$-separation statement in the acyclified graph $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$. Consequently, the acyclified graph $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$ is equivalent to the augmented graph $\mathcal{G}^{\mathcal{I}}$. Furthermore, by the results of [51], $p_{\mathcal{G}^{\mathcal{I}}}$ admits a factorization, as formalized in the theorem below.
Theorem A.11 ([51]).
@@
From the above definition it is clear that $\mathcal{P}_{\mathcal{I}}(\mathcal{G})\subseteq\mathcal{M}_{\mathcal{I}}(\mathcal{G})$. Theorem 3.1 relies on the following set of assumptions. The first one ensures that the model is capable of representing the ground truth distribution.
@@
The above assumption ensures that the hypothetical scenario where $\mathcal{S}(\mathcal{G}^{\ast})$ and $\mathcal{S}(\mathcal{G})$ are both infinity is avoided. This is formalized in the lemma below taken from [15].
Lemma A.16 (Finiteness of the score function [15]).
@@
From the results of [15], we can now express the difference in score function between $\mathcal{G}^{\ast}$ and $\mathcal{G}$ as the minimization of KL diverengence plus the difference in the regularization terms.
Lemma A.17 (Rewritting the score function [15]).
@@
We will now prove the following technical lemma (adapted from [15]) which we will be used in proving Theorem 3.1.
@@
We are now ready to prove Theorem 3.1. Recall,
Theorem 3.1.
Let $\mathcal{I}=\{I_{k}\}_{k=0}^{K}$ be a family of interventional targets, let $\mathcal{G}^{\ast}$ denote the ground truth directed mixed graph, $p^{(k)}$ denote the data generating distribution for $I_{k}$, and $\hat{\mathcal{G}}:=\arg\max_{\mathcal{G}}\mathcal{S}(\mathcal{G})$. Then, under the Assumptions 2.1, A.13, A.14, and A.15, and for a suitably chosen $\lambda>0$, we have that $\hat{\mathcal{G}}\equiv_{\mathcal{I}}\mathcal{G}^{\ast}$. That is, $\hat{\mathcal{G}}$ is $\mathcal{I}$-Markov equivalent to $\mathcal{G}^{\ast}$.
@@
It is sufficient to show that for $\mathcal{G}\notin\mathcal{I}\text{-MEC}(\mathcal{G}^{\ast})$, the score function of $\hat{\mathcal{G}}$ is strictly lower than the score function of $\mathcal{G}^{\ast}$, i.e., $\mathcal{S}(\mathcal{G}^{\ast})>\mathcal{S}(\mathcal{G})$. Since $\mathcal{G}\notin\mathcal{I}\text{-MEC}(\mathcal{G}^{\ast})$ and $p_{\mathcal{I}}^{\ast}\in\mathcal{M}_{\mathcal{I}}(\mathcal{G}^{\ast})$ (by Assumption A.13), there must exist subsets of nodes $A,B,C\subseteq\mathcal{V}\cup\bm{C}^{\mathcal{I}}$ such that
$A\mathop{\perp}_{\mathcal{G}}^{\sigma}B\mid C\quad\text{and}\quad A\mathop{\not\perp}_{\mathcal{G}^{\ast}}^{\sigma}B\mid C.$
@@
Following [15], we now show that by choosing $\lambda$ sufficiently small, the above equation is stictly positive. Note that if $|\mathcal{G}|\geq|\mathcal{G}^{\ast}|$ then $\mathcal{S}(\mathcal{G}^{\ast})-\mathcal{S}(\mathcal{G})>0$. Let $\mathbb{G}^{+}:=\{\mathcal{G}\mid|\mathcal{G}|<|\mathcal{G}^{\ast}|\}$. Choosing $\lambda$ such that $0<\lambda<\min_{\mathcal{G}\in\mathbb{G}^{+}}\frac{\eta(\mathcal{G})}{|\mathcal{G}^{\ast}|-|\mathcal{G}|}$ we see that:
@@ A.4 Characterization of Equivalence Class
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{B})$ be a directed mixed graph, and consider a family of interventional targets $\mathcal{I}={I_{k}}_{k=0}^{K}$ with $I_{0}=\emptyset$. From Proposition A.6, for any graph $\mathcal{G}_{1}\in\mathcal{I}\text{-MEC}(\mathcal{G})$, the corresponding augmented graph $\mathcal{G}_{1}^{\mathcal{I}}$ is equivalent to the acyclification $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$ of $\mathcal{G}^{\mathcal{I}}$. Several prior works have studied the characterization of equivalence classes of acyclic directed mixed graphs (ADMGs), including [53, 30, 54]. We now provide a graphical notion of the $\mathcal{I}$-Markov equivalence class of a DMG $\mathcal{G}$. A graph $\mathcal{G}$ is said to be maximal if there exists no inducing path (relative to the empty set) between any two non-adjacent nodes. An inducing path relative to a subset $L$ is a path on which every non-endpoint node $i\notin L$ is a collider on the path and every collider is an ancestor of an endpoint of the path. A Maximal Ancestral Graph (MAG) is one that is both ancestral and maximal. Given an ADMG $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$, it is possible to construct a MAG over the variable set $\bm{V}=(\bm{X},\bm{C}^{\mathcal{I}})$ that preserves both the independence structure and ancestral relationships encoded in $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$; see [55] for details. We denote $MAG(\mathrm{acy}(\mathcal{G}^{\mathcal{I}}))$ to mean MAG that is constructed from $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$. Therefore all the independencies encoded in $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$ is also present in $MAG(\mathrm{acy}(\mathcal{G}^{\mathcal{I}}))$. Before present the condition for MAG equivalence, we introduce the notion of discriminating path. A path $\pi=(i_{0},\varepsilon_{1},\ldots,i_{n-1},\varepsilon_{n},i_{n})$ in $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$ is called a discriminating path for $i_{n-1}$ if (1) $\pi$ includes at lest three edges; (2) $i_{n-1}$ is a non-endpoint node on $\pi$, and is adjacent to $i_{n}$ on $\pi$; and (3) $i_{0}$ and $i_{n}$ are not adjacent, and every node in between $i_{0}$ and $i_{n-1}$ is a collider on $\pi$ and is a parent of $i_{n}$. The following theorem from Spirtes and Richardson [53] characterizes the equivalence of MAGs.
@@
Therefore, from Proposition A.6 and Theorem A.19, two DMGs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are equivalent if and only if $MAG(\mathrm{acy}(\mathcal{G}_{1}^{\mathcal{I}}))$ and $MAG(\mathrm{acy}(\mathcal{G}_{2}^{\mathcal{I}}))$ satisfying the conditions of Theorem 3.1, i.e., $MAG(\mathrm{acy}(\mathcal{G}_{1}^{\mathcal{I}}))$ and $MAG(\mathrm{acy}(\mathcal{G}_{2}^{\mathcal{I}}))$: (i) have the same skeleton, (ii) same unshielded colliders, and (iii) same discriminating paths with consistent colliders.

### v2

@@ 80. Guidelines:
The appendices are structured as follows: Appendix A presents the theoretical foundations of differentiable cyclic causal discovery in the presence of unmeasured confounders, including the proof of Theorem 2, and a characterization of the equivalence class of DMGs that maximize the score function. Appendix B provides implementation details of DCCD-CONF and the baselines. Finally, Appendix C provides additional experimental results comparing DCCD-CONF with the baselines.
@@ Appendix A Theory
In this section, we establish the theoretical foundations of differentiable cyclic causal discovery in the presence of unmeasured confounders. We begin by reviewing key definitions and results from prior work that are essential for proving Theorem 2, starting with fundamental graph terminology.
@@ A.1 Preliminaries
where $\mathop{\perp}_{p_{\mathcal{G}}}$ denotes conditional independence, and $p_{\mathcal{G}}$ denotes the observational distribution. This is known as the directed global Markov property of $\mathcal{G}$ [Forré and Mooij, 2017]. However, in general, cyclic graphs do not obey the directed global Markov property as shown by the counterexample below taken from [Bongers et al., 2021, Spirtes, 2013].
@@
Forré and Mooij [2017] introduced $\sigma$-separation as a generalization of d-separation to extend the directed global Markov property to cyclic graphs. This concept was motivated by applying $d$-separation to the acyclified version of the DMG. Before delving into $\sigma$-separation, we first define the acyclification procedure of a directed mixed graph, following [Bongers et al., 2021].
@@
Note that $\sigma$-separation reduces to $d$-separation for acyclic graphs, that is, when $\mathrm{sc}_{\mathcal{G}}(i)=\{i\}$ for all $i\in\mathcal{V}$. The following result in [Forré and Mooij, 2017] relates $\sigma$-separation and $d$-separation.
Proposition A.6 (Forré and Mooij [2017]).
@@
Definition A.7 (General directed global Markov property [Forré and Mooij, 2017]).
@@ A.2 Joint Causal Modelling and Markov properties
In order to incorporate multiple interventional settings into a single causal modeling framework, we follow the joint causal model introduced by Mooij et al. [2020], where we augment the system with a set of context variables $\bm{C}^{\mathcal{I}}=(\bm{C}_{1},\ldots,\bm{C}_{K})$ each corresponding to a non-empty interventional setting. In this case, $\bm{C}_{k}=\emptyset$ for all $k=1,\dots,K$ corresponds to the observational setting. We construct an augmented graph, denoted by $\mathcal{G}^{\mathcal{I}}$ consisting of both the system variables $\bm{X}$ and the context variables $\bm{C}^{\mathcal{I}}$, such that the $\mathrm{ch}_{\mathcal{G}}(\bm{C}_{k})=I_{k}$, and no context variable has any parent or a spouse. Figure 5 illustrates the augmented graph for the graph from Example A.3 and the intervention sets $\mathcal{I}=\{\emptyset,\{X_{3}\},\{X_{4}\}\}$. The new system containing both the observed variables and the context variables is called the meta system. Finally, given a family of interventional targets $\mathcal{I}=\{I_{k}\}_{k=1}^{K}$ and the corresponding context variable $\bm{C}_{k}$, the structural equations of the meta systems governing the observations $\bm{X}$ and $\bm{C}^{\mathcal{I}}$ has the following form:
$\tilde{F}_{i}(\bm{X}_{\mathrm{pa}_{\mathcal{G}}(i)},\bm{C}^{\mathcal{I}}_{\mathrm{pa}_{\mathcal{G}}^{\mathcal{I}}(i)},Z_{i})=\begin{cases}(\bm{C}_{k})_{i},&\text{if }\exists\,k\in[K]\text{ s.t. }\bm{C}_{K}\neq\emptyset,\text{ and }X_{i}\in I_{k},\\ F_{i}(\bm{X}_{\mathrm{pa}_{\mathcal{G}}(i)},Z_{i}),&\text{otherwise}.\end{cases}$
@@
For an DMG $\mathcal{G}$ and a choice of $\mathbf{F}:\mathbb{R}^{2d}\to\mathbb{R}^{d}$ such that $\mathbf{f}_{x}$ is unique and invertible and $\mathbf{\Sigma}\succ 0$, the structural equations are uniquely solvable with respect to each strongly connected component of $\mathcal{G}$. Morover, the addition of context variables in the augmented graph does not introduce any new cycles. Therefore the meta system forms a simple SCM. Thus, from Theorem A.21 in [Bongers et al., 2021], the distribution $p_{\mathcal{G}^{\mathcal{I}}}$ is unique and it satisfies the general directed global Markov property. ∎
@@
From Proposition A.6, for a DMG $\mathcal{G}$ and a family of interventional targets $\mathcal{I}=\{I_{k}\}_{k=0}^{K}$, any $\sigma$-separation statement in $\mathcal{G}^{\mathcal{I}}$ translates to a $d$-separation statement in the acyclified graph $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$. Consequently, the acyclified graph $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$ is equivalent to the augmented graph $\mathcal{G}^{\mathcal{I}}$. Furthermore, by the results of [Richardson, 2009], $p_{\mathcal{G}^{\mathcal{I}}}$ admits a factorization, as formalized in the theorem below.
Theorem A.11 (Richardson [2009]).
@@ A.3 Proof of Theorem 2
From the above definition it is clear that $\mathcal{P}_{\mathcal{I}}(\mathcal{G})\subseteq\mathcal{M}_{\mathcal{I}}(\mathcal{G})$. Theorem 2 relies on the following set of assumptions. The first one ensures that the model is capable of representing the ground truth distribution.
@@
The above assumption ensures that the hypothetical scenario where $\mathcal{S}(\mathcal{G}^{\ast})$ and $\mathcal{S}(\mathcal{G})$ are both infinity is avoided. This is formalized in the lemma below taken from [Brouillard et al., 2020].
Lemma A.16 (Finiteness of the score function [Brouillard et al., 2020]).
@@
From the results of [Brouillard et al., 2020], we can now express the difference in score function between $\mathcal{G}^{\ast}$ and $\mathcal{G}$ as the minimization of KL diverengence plus the difference in the regularization terms.
Lemma A.17 (Rewritting the score function [Brouillard et al., 2020]).
@@
We will now prove the following technical lemma (adapted from [Brouillard et al., 2020]) which we will be used in proving Theorem 2.
@@
This proves the lemma. ∎
We are now ready to prove Theorem 2. Recall,
Theorem 2.
Let $\mathcal{I}=\{I_{k}\}_{k=0}^{K}$ be a family of interventional targets, let $\mathcal{G}^{\ast}$ denote the ground truth directed mixed graph, $p^{(k)}$ denote the data generating distribution for $I_{k}$, and $\hat{\mathcal{G}}:=\arg\max_{\mathcal{G}}\mathcal{S}(\mathcal{G})$. Then, under the Assumptions 1, A.13, A.14, and A.15, and for a suitably chosen $\lambda>0$, we have that $\hat{\mathcal{G}}\equiv_{\mathcal{I}}\mathcal{G}^{\ast}$. That is, $\hat{\mathcal{G}}$ is $\mathcal{I}$-Markov equivalent to $\mathcal{G}^{\ast}$.
@@
It is sufficient to show that for $\mathcal{G}\notin\mathcal{I}\text{-MEC}(\mathcal{G}^{\ast})$, the score function of $\hat{\mathcal{G}}$ is strictly lower than the score function of $\mathcal{G}^{\ast}$, i.e., $\mathcal{S}(\mathcal{G}^{\ast})>\mathcal{S}(\mathcal{G})$. Since $\mathcal{G}\notin\mathcal{I}\text{-MEC}(\mathcal{G}^{\ast})$ and $p_{\mathcal{I}}^{\ast}\in\mathcal{M}_{\mathcal{I}}(\mathcal{G}^{\ast})$ (by Assumption A.13), there must exist subsets of nodes $A,B,C\subseteq\mathcal{V}\cup\bm{C}^{\mathcal{I}}$ such that either:
$A\mathop{\perp}_{\mathcal{G}}^{\sigma}B\mid C\quad\text{and}\quad A\mathop{\not\perp}_{\mathcal{G}^{\ast}}^{\sigma}B\mid C,$
or
$A\mathop{\not\perp}_{\mathcal{G}}^{\sigma}B\mid C\quad\text{and}\quad A\mathop{\perp}_{\mathcal{G}^{\ast}}^{\sigma}B\mid C,$
@@
Following [Brouillard et al., 2020], we now show that by choosing $\lambda$ sufficiently small, the above equation is stictly positive. Note that if $|\mathcal{G}|\geq|\mathcal{G}^{\ast}|$ then $\mathcal{S}(\mathcal{G}^{\ast})-\mathcal{S}(\mathcal{G})>0$. Let $\mathbb{G}^{+}:=\{\mathcal{G}\mid|\mathcal{G}|<|\mathcal{G}^{\ast}|\}$. Choosing $\lambda$ such that $0<\lambda<\min_{\mathcal{G}\in\mathbb{G}^{+}}\frac{\eta(\mathcal{G})}{|\mathcal{G}^{\ast}|-|\mathcal{G}|}$ we see that:
@@ A.4 Characterization of Equivalence Class
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{B})$ be a directed mixed graph, and consider a family of interventional targets $\mathcal{I}={I_{k}}_{k=0}^{K}$ with $I_{0}=\emptyset$. From Proposition A.6, for any graph $\mathcal{G}_{1}\in\mathcal{I}\text{-MEC}(\mathcal{G})$, the corresponding augmented graph $\mathcal{G}_{1}^{\mathcal{I}}$ is equivalent to the acyclification $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$ of $\mathcal{G}^{\mathcal{I}}$. Several prior works have studied the characterization of equivalence classes of acyclic directed mixed graphs (ADMGs), including Spirtes and Richardson [1997], Kocaoglu et al. [2019], Ali et al. [2009]. We now provide a graphical notion of the $\mathcal{I}$-Markov equivalence class of a DMG $\mathcal{G}$. A graph $\mathcal{G}$ is said to be maximal if there exists no inducing path (relative to the empty set) between any two non-adjacent nodes. An inducing path relative to a subset $L$ is a path on which every non-endpoint node $i\notin L$ is a collider on the path and every collider is an ancestor of an endpoint of the path. A Maximal Ancestral Graph (MAG) is one that is both ancestral and maximal. Given an ADMG $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$, it is possible to construct a MAG over the variable set $\bm{V}=(\bm{X},\bm{C}^{\mathcal{I}})$ that preserves both the independence structure and ancestral relationships encoded in $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$; see Zhang [2008] for details. We denote $MAG(\mathrm{acy}(\mathcal{G}^{\mathcal{I}}))$ to mean MAG that is constructed from $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$. Therefore all the independencies encoded in $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$ is also present in $MAG(\mathrm{acy}(\mathcal{G}^{\mathcal{I}}))$. Before present the condition for MAG equivalence, we introduce the notion of discriminating path. A path $\pi=(i_{0},\varepsilon_{1},\ldots,i_{n-1},\varepsilon_{n},i_{n})$ in $\mathrm{acy}(\mathcal{G}^{\mathcal{I}})$ is called a discriminating path for $i_{n-1}$ if (1) $\pi$ includes at lest three edges; (2) $i_{n-1}$ is a non-endpoint node on $\pi$, and is adjacent to $i_{n}$ on $\pi$; and (3) $i_{0}$ and $i_{n}$ are not adjacent, and every node in between $i_{0}$ and $i_{n-1}$ is a collider on $\pi$ and is a parent of $i_{n}$. The following theorem from Spirtes and Richardson [1997] characterizes the equivalence of MAGs.
@@
Therefore, from Proposition A.6 and Theorem A.19, two DMGs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are equivalent if and only if $MAG(\mathrm{acy}(\mathcal{G}_{1}^{\mathcal{I}}))$ and $MAG(\mathrm{acy}(\mathcal{G}_{2}^{\mathcal{I}}))$ satisfying the conditions of Theorem 2, i.e., $MAG(\mathrm{acy}(\mathcal{G}_{1}^{\mathcal{I}}))$ and $MAG(\mathrm{acy}(\mathcal{G}_{2}^{\mathcal{I}}))$: (i) have the same skeleton, (ii) same unshielded colliders, and (iii) same discriminating paths with consistent colliders.